OpenStudy (anonymous):

7a - b = 15a solve for a

OpenStudy (jack1):

so get everything with an "a" in it on one side and everything with a b in it on the other side of the equals sign

OpenStudy (anonymous):

7a-b=15a -b=8a a=-b/8

OpenStudy (jack1):

yep, perfect

OpenStudy (anonymous):

thanks